OpenStudy (anonymous):

what is the value of 8/36 in lowest terms? A.4/18 B.1/2 C. 2/9 D. 1/4

OpenStudy (zzr0ck3r):

divide the top and bottom by 4

OpenStudy (anonymous):

2/9

OpenStudy (anonymous):

jk its 1/4

OpenStudy (zzr0ck3r):

@cantthink96 you should explain your answers. dont just give answers... this helps no one

OpenStudy (zzr0ck3r):

its not \(\frac{1}{4}\)

OpenStudy (anonymous):

let me explain lad, when finding lowest term you want to find whats the largest number that goes into the top and bottom, i correct myself it is not 1/4 and i got back to my 2/9 statment because it is not 32 but 36

OpenStudy (zzr0ck3r):

\[\frac{8}{36}=\frac{\frac{8}{4}}{\frac{36}{4}}=\frac{2}{9}\]